OpenStudy (anonymous):

What are the restrictions on -1/6 = 9/3x

OpenStudy (caominhim):

well i think the most constraining restriction would be the fact that there is only 1 possible x here

OpenStudy (anonymous):

So it would be 3 or do you divide by 3

OpenStudy (caominhim):

\[\frac{-\frac{1}{6}}{\frac{9}{3}} =x\]

OpenStudy (jhannybean):

\[\huge \frac{\dfrac{a}{b}}{\dfrac{c}{d}} = \frac{a}{b} \cdot \frac{d}{c}\]